# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v2

@@ 5.2. Proof of Theorems 1 and 2
The assumptions (iii) and (iv) hold by Proposition 7. To see that this proposition is applicable, it only remains to check that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. Fix $\rho\in\{+,-\}$. Since $s-g_{\rho}(s)=f_{\rho}(g_{\rho}(s))$ for $s\geq f_{\rho}(0)$ and $g_{\rho}(s)\to+\infty$ as $s\to+\infty$, the integral $\int_{f_{\rho}(0)}^{+\infty}|g_{\rho}(s)-s|\,ds$ is finite only if $f_{\rho}(+\infty)=0$. At the same time, (15) implies that this integral is finite almost surely, hence $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ holds as claimed. Thus, the hypotheses of Proposition 7 are satisfied, and it follows that assumptions (iii) and (iv) both apply to the sequence $\{T^{(N)}\}$.
@@
According to Proposition 11 above, from the fact that $\sigma\neq 0$, assumption (vii) also holds. Therefore, Corollary 4 applies, establishing convergence of the rescaled CDF $F_{N}$ as desired to the solution of the initial-value problem associated to the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0$, as desired.
@@ 5.3. Proof of Corollary 1
$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}N^{1/2}}X^{(N)}\leq x\}=F(2x\sqrt{|\sigma|},1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
The righthand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Therefore, $2^{-1}N^{-1/2}X^{(N)}\overset{d}{\to}\sqrt{|\sigma|}\text{Beta}(2,1)$.
@@
This is equivalent to asking that $Y^{(N,0)}=X^{(0)}+N^{1/2}Z^{(N)}$ for some sequence of nonnegative random variables $\{Z^{(N)}\}$ with law converging to $G_{\text{in}}$.
@@
and thus $X^{(n)}\overset{d}{=}(1-\Theta)\max\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}+\Theta\min\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}\overset{d}{=}X^{(n-1)}$ for all $n$ (since $\mathbb{P}\{\Theta=1\}=p=1/2$). In other words, $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$, completing the proof of case (iii). ∎
@@ 5.4. Proof of Proposition 10
With these preliminaries out of the way, fix a $\Phi\in CDF(\overline{\mathbb{R}})$ that has bounded, uniformly continuous first and second derivatives, and define $\Phi_{\delta}(y)=\Phi(\delta y)$ for $\delta>0$. Let $\omega_{\Phi}$ be the modulus of continuity
$\omega_{\Phi}(\delta)=\sup\left\{|\partial_{x}\Phi(x)-\partial_{x}\Phi(y)|+|\partial_{x}^{2}\Phi(x)-\partial_{x}^{2}\Phi(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}.$
@@
for some $N$-independent functions $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$ defined on $CDF(\overline{\mathbb{R}})$. Thus, the scaling behavior of $\mathscr{L}^{(N)}$ is determined by that of $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$. Further, since the assumption (15) implies that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=1\}$ (as proved in Section 5.2 above), the formula (33) can be invoked to find
$(\mathscr{L}_{+}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x-\delta h_{+}(s))-\Phi(x)]\partial_{x}\Phi(x-\delta s)\,ds,$
$(\mathscr{L}_{-}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x+\delta h_{-}(s))-\Phi(x)]\partial_{x}\Phi(x+\delta s)\,ds.$
@@
(56) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
@@
(57) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
@@
(58) $\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}$
$=\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds.$
By the assumptions on $\Phi$ and the bound $h_{\rho}(s)\leq s$,
$\left|\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds\right|$
$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\mathbf{E}\int_{0}^{+\infty}\left(\omega_{\Phi}(\delta h_{+}(s))+\omega_{\Phi}(\delta s)\right)h_{\rho}(s)\,ds$
$\qquad\leq 2\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)\,ds$
@@
$\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds$
$\quad=\rho\partial_{x}\Phi(x)\mathbf{E}\int_{0}^{+\infty}(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds$
$\quad\quad+\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds.$
@@
$\left|\frac{\rho}{\delta}\mathbf{E}\int_{0}^{+\infty}\partial_{x}\Phi(x)(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds-\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}s\mathbf{E}h_{\rho}(s)\,ds\right|$
$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
Similarly, in the second term, again using $\omega_{\Phi}(\delta h_{+}(s))\leq\omega_{\Phi}(\delta s)$,
$\left|\frac{1}{\delta}\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds\right$.
$\qquad\left.-\frac{1}{2}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}\mathbf{E}h_{\rho}(s)^{2}\,ds\right|$
$\qquad\leq\frac{1}{2}\left(\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)^{2}\,ds.$
@@
$\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}$
$\qquad\leq\left(1+\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
@@ 5.5. Counterexample
where $Z$ is a nonnegative random variable such that $\mathbb{E}Z^{2}=+\infty$. With this choice of forcing, the RDE (9) with $p=1/2$ is superdiffusive in the sense that if $\{X^{(n)}\}$ is a solution started from any initial datum and $p=1/2$, then
@@
Let $\{X^{(m,n)}\}$ be a solution of the RDE (9) started from the same initial distribution as $\{X^{(n)}\}$, but with the forcing $(f_{+},0)$ replaced by $(f_{+}^{(m)},0)$. Note that the inverse $g_{+}^{(m)}$ associated with $f_{+}^{(m)}$ is given by
@@
and the inverse $g_{-}$ associated with $f_{-}=0$ is $g_{-}(s)=s$. Therefore, the RDE with $(f^{(m)}_{+},0)$ satisfies (15) and the constant $\sigma^{(m)}$ determining the asymptotics is given by
@@
$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}\leq F(2^{-1}|\sigma^{(m)}|^{-\frac{1}{2}}x).$
where $F$ is the CDF of the $\text{Beta}(2,1)$ random variable Sending $m\to+\infty$, this implies
$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq F(0)=0\quad\text{for each}\,\,x\in\mathbb{R}.$
@@ Appendix A Viscosity Solutions
(64) $\mathcal{F}(0,0)=0,\quad Q(0)\geq 0,\quad\text{and}\quad Q(1)\leq 0.$
@@
An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
@@
A lower semi-continuous function $G:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)+Q(G)\geq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
@@
A continuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0$ in the viscosity sense if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\geq 0$ in the viscosity sense.
@@
An upper semicontinuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ said to be a viscosity subsolution of (61) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
@@
A lower semicontinuous function $G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (61) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
@@
A continuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a (continuous) viscosity solution of (61) if it is both a viscosity subsolution and a viscosity supersolution.
Since the main results of this work, Theorems 1 and 2, involve problems with a discontinuous initial datum, a different definition of solution is needed. In order to state it, it is necessary to recall the notions of upper and lower semicontinuous envelopes. Specifically, for a function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$, the upper and lower semicontinuous envelopes $F^{*}$ and $F_{*}$ are the functions defined in $\mathbb{R}\times[0,\infty)$ via the formulae
@@
A function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (61) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})+Q(F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})+Q(F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; and $F^{*}(\cdot,0)\leq(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)\geq(F_{\text{in}})_{*}$ in $\mathbb{R}$.

### v3

@@ 5.2. Proof of Theorems 1 and 2
The assumptions (iii) and (iv) hold by Proposition 7. To see that this proposition is applicable, it only remains to check that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. Fix $\rho\in\{+,-\}$. Since $s-g_{\rho}(s)=f_{\rho}(g_{\rho}(s))$ for $s\geq f_{\rho}(0)$ and $g_{\rho}(s)\to+\infty$ as $s\to+\infty$, the integral $\int_{f_{\rho}(0)}^{+\infty}|g_{\rho}(s)-s|\,ds$ is finite only if $f_{\rho}(+\infty)=0$. At the same time, (15) implies that this integral is finite almost surely, hence $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ as claimed. Thus, the hypotheses of Proposition 7 are satisfied, and it follows that assumptions (iii) and (iv) both apply to the sequence $\{T^{(N)}\}$.
@@
According to Proposition 11 above, from the fact that $\sigma\neq 0$, assumption (vii) also holds. Therefore, Corollary 4 applies, establishing convergence of the rescaled CDF $F_{N}$ to the solution of the initial-value problem associated to the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0$, as desired.
@@
Remark 3.
Notice that, in the proof of Theorems 1 and 2 above, nowhere was it necessary to assume that $\{X^{(N,n)}\}$ was real-valued. Therefore, since Theorem 3 and Corollary 4 in particular do not require that $F^{(N)}_{n}(+\infty)=1$ or $F^{(N)}_{n}(-\infty)=0$, the conclusions of the theorems remain true if $\{X^{(N,n)}\}$ are extended real-valued.
@@ 5.3. Proof of Corollary 1
$\lim_{N\to\infty}\mathbb{P}\{-\text{sgn}(\sigma)(2\sqrt{|\sigma|})^{-1}N^{-\frac{1}{2}}X^{(N)}\leq x\}=F(2x\sqrt{|\sigma|},1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
The righthand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Therefore, $2^{-1}N^{-1/2}X^{(N)}\overset{d}{\to}-\text{sgn}(\sigma)\sqrt{|\sigma|}\text{Beta}(2,1)$.
@@
This is equivalent to asking that $Y^{(N,0)}\overset{d}{=}X^{(0)}+N^{1/2}Z^{(N)}$ for some sequence of nonnegative random variables $\{Z^{(N)}\}$ with law converging to $G_{\text{in}}$.
@@
and thus $X^{(n)}\overset{d}{=}(1-\Theta)\max\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}+\Theta\min\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}\overset{d}{=}X^{(n-1)}$ for all $n$ (since $\mathbb{P}\{\Theta=1\}=p=1/2$). In other words, $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$, completing the proof in case (iii). ∎
@@ 5.4. Proof of Proposition 10
With these preliminaries out of the way, fix a $G\in CDF(\overline{\mathbb{R}})$ that has bounded, uniformly continuous first and second derivatives, and define $G_{\delta}(y)=G(\delta y)$ for $\delta>0$. Let $\omega_{G}$ be the modulus of continuity
$\omega_{G}(\delta)=\sup\left\{|\partial_{x}G(x)-\partial_{x}G(y)|+|\partial_{x}^{2}G(x)-\partial_{x}^{2}G(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}.$
@@
for some $N$-independent functions $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$ defined on $CDF(\overline{\mathbb{R}})$. Thus, the scaling behavior of $\mathscr{L}^{(N)}$ is determined by that of $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$. Further, since the assumption (15) implies that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ (as proved in Section 5.2 above), the formula (33) can be invoked to find
$(\mathscr{L}_{+}G_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[G(x-\delta h_{+}(s))-G(x)]\partial_{x}G(x-\delta s)\,ds,$
$(\mathscr{L}_{-}G_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[G(x+\delta h_{-}(s))-G(x)]\partial_{x}G(x+\delta s)\,ds.$
@@
(56) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-2}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}G(x)|^{2}\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
@@
(57) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}G(x)|^{2}-a_{\rho}|\partial_{x}G(x)|\partial_{x}^{2}G(x)\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
@@
(58) $\delta^{-2}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}G(x)|^{2}$
$=\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}\partial_{x}G(x-\rho\delta s)+\rho|\partial_{x}G(x)|^{2}\right)h_{\rho}(s)\,ds.$
By the assumptions on $G$ and the bound $h_{\rho}(s)\leq s$,
$\left|\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}\partial_{x}G(x-\rho\delta s)+\rho|\partial_{x}G(x)|^{2}\right)h_{\rho}(s)\,ds\right|$
$\qquad\leq\|\partial_{x}G\|_{\text{sup}}\mathbf{E}\int_{0}^{+\infty}\left(\omega_{G}(\delta h_{\rho}(s))+\omega_{G}(\delta s)\right)h_{\rho}(s)\,ds$
$\qquad\leq 2\|\partial_{x}G\|_{\text{sup}}\int_{0}^{+\infty}\omega_{G}(\delta s)\mathbf{E}h_{\rho}(s)\,ds$
@@
$\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}\partial_{x}G(x-\rho\delta s)+\rho|\partial_{x}G(x)|^{2}\right)h_{\rho}(s)\,ds$
$\quad=\rho\partial_{x}G(x)\mathbf{E}\int_{0}^{+\infty}(\partial_{x}G(x)-\partial_{x}G(x-\rho\delta s))h_{\rho}(s)\,ds$
$\quad\quad+\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}+\rho\partial_{x}G(x)\right)\partial_{x}G(x-\rho\delta s)h_{\rho}(s)\,ds.$
@@
$\left|\frac{\rho}{\delta}\mathbf{E}\int_{0}^{+\infty}\partial_{x}G(x)(\partial_{x}G(x)-\partial_{x}G(x-\rho\delta s))h_{\rho}(s)\,ds-\partial_{x}G(x)\partial_{x}^{2}G(x)\int_{0}^{+\infty}s\mathbf{E}h_{\rho}(s)\,ds\right|$
$\qquad\leq\|\partial_{x}G\|_{\text{sup}}\int_{0}^{+\infty}\omega_{G}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
Similarly, in the second term, again using $\omega_{G}(\delta h_{\rho}(s))\leq\omega_{G}(\delta s)$,
$\left|\frac{1}{\delta}\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}+\rho\partial_{x}G(x)\right)\partial_{x}G(x-\rho\delta s)h_{\rho}(s)\,ds\right$.
$\qquad\left.-\frac{1}{2}\partial_{x}G(x)\partial_{x}^{2}G(x)\int_{0}^{+\infty}\mathbf{E}h_{\rho}(s)^{2}\,ds\right|$
$\qquad\leq\frac{1}{2}\left(\|\partial_{x}G\|_{\text{sup}}+\|\partial_{x}^{2}G\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{G}(\delta s)\mathbf{E}h_{\rho}(s)^{2}\,ds.$
@@
$\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}G(x)|^{2}-a_{\rho}|\partial_{x}G(x)|\partial_{x}^{2}G(x)\right|\,\mid\,x\in\mathbb{R}\right\}$
$\qquad\leq\left(1+\|\partial_{x}G\|_{\text{sup}}+\|\partial_{x}^{2}G\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{G}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
@@ 5.5. Counterexample
where $Z$ is a nonnegative random variable such that $\mathbb{E}Z^{2}=+\infty$. With this choice of forcing, the RDE (9) with $p=1/2$ is superdiffusive in the sense that if $\{X^{(n)}\}$ is a solution started from any real-valued initial datum, then
@@
Let $\{X^{(m,n)}\}$ be a solution of the RDE (9) started from the same initial distribution as $\{X^{(n)}\}$, but with the forcing $(f_{+},0)$ replaced by $(f_{+}^{(m)},0)$. Note that the function $g_{+}^{(m)}$ associated to $f_{+}^{(m)}$ via (14) is given by
@@
and the function $g_{-}$ associated to $f_{-}=0$ is $g_{-}(s)=s$. Therefore, the RDE with $(f^{(m)}_{+},0)$ satisfies (15) and the constant $\sigma^{(m)}$ determining the asymptotics is given by
@@
$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}\leq F_{\text{Beta}(2,1)}(2^{-1}|\sigma^{(m)}|^{-\frac{1}{2}}x).$
where $F_{\text{Beta}(2,1)}$ is the CDF of the $\text{Beta}(2,1)$ random variable. Sending $m\to+\infty$, this implies
$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq F_{\text{Beta}(2,1)}(0)=0\quad\text{for each}\,\,x\in\mathbb{R}.$
@@ Appendix A Viscosity Solutions
(64) $\mathcal{F}(0,0)=0,\quad Q(0)\leq 0,\quad\text{and}\quad Q(1)\geq 0.$
@@
An upper semicontinuous function $F:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ if it has the following property: If $\varphi$ is a smooth function defined in some open set $U\subseteq\mathbb{R}\times(0,+\infty)$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
@@
A lower semicontinuous function $G:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ if it has the following property: If $\varphi$ is a smooth function defined in some open set $U\subseteq\mathbb{R}\times(0,+\infty)$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
@@
A continuous function $F:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$.
@@
An upper semicontinuous function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a viscosity subsolution of (61) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and
@@
A lower semicontinuous function $G:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (61) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and
@@
A continuous function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a (continuous) viscosity solution of (61) if it is both a viscosity subsolution and a viscosity supersolution.
Since the main results of this work, Theorems 1 and 2, consider the problem (61) with initial data that may be discontinuous, a different definition of solution is needed. In order to state it, it is necessary to recall the notions of upper and lower semicontinuous envelopes. Specifically, for a function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$, the upper and lower semicontinuous envelopes $F^{*}$ and $F_{*}$ are the functions defined in $\mathbb{R}\times[0,+\infty)$ via the formulae
@@
A function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (61) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})+Q(F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})+Q(F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$; and $F^{*}(\cdot,0)\leq(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)\geq(F_{\text{in}})_{*}$ in $\mathbb{R}$.